OpenStudy (anonymous):

Can someone walk me through this!? f(x) = -2x+4 and g(x)= -6x-7 find f(x)-g(x). I know the answer is 4-2x^2 I just don't know the steps to get that answer.

OpenStudy (mathstudent55):

That can't be the solution if your functions and operation are correct. Neither function has an x^2 term. Subtraction can't result in an x^2 term.

OpenStudy (anonymous):

Can you help explain it then? I'm really stuck on this.

OpenStudy (mathstudent55):

I can help you with the problem you posted, but you're not going to get the answer you wrote.

OpenStudy (mathstudent55):

\(\color{red}{f(x) = -2x+4}\) \(\color{green}{g(x)= -6x-7}\) \(\color{red}{f(x)}-\color{green}{g(x)} \) \(~~~~~~~~~~~~~~~~~= \color{red}{-2x + 4} - \color{green}{(-6x-7)}\) \(~~~~~~~~~~~~~~~~~= -2x + 4 + 6x + 7\) \(f(x) - g(x) = 4x + 11\)

OpenStudy (anonymous):

Wow thats actually really simple thank you so much! Could you help me with another problem kind of similar? I can try solving it and if i'm wrong could you correct me?

OpenStudy (mathstudent55):

I'm getting off in a few minutes. Post it anyway if you need help. If I'm not here, someone else will help you.

OpenStudy (anonymous):

okay thanks for the help though!